OpenStudy (anonymous):

Numerical Methods question. When calculating an number N, whe get an approximation of 29,25, with a error percentage of <10%. Determine the limits in wich N should be contained. Sol: N ϵ [26.591, 31.909]. How do I get there?

OpenStudy (ikram002p):

29,25 + 10% of 29,25 or 29,25 - 10% of 29,25 but im getting [26.325,32.175] so sryyy mmm :-'(

OpenStudy (anonymous):

Yes, that's what I get, and that's the answer for the second part of the question.

OpenStudy (ikram002p):

what is the second part of the question ?

OpenStudy (anonymous):

solve the previous question again but assuming that the approximation is made by defect* Sol: (29.25; 32.175) *defect is probably not the english term for what I'm asking, but I don't know how this kind of approximation is called.

OpenStudy (mathmate):

29,25 is either over or underestimated by 10%. In the first case, the accurate value would be 29,25/1,10=26,591 (26,591*1,1=29,25) In the second case, the accurate value is 29.25/0.9=32,5, different from the answer as well. (32,5*0.9=29,25) By the way, what was "defect" in its original language?

OpenStudy (anonymous):

an approximation by defect - uma aproximação por defeito